# On Doubly Robust Estimation with Nonignorable Missing Data Using Instrumental Variables

### v2

@@ 3.2 Influence function-based doubly robust estimators
Following Robins and Rotnitzky, (2001) and Chernozhukov et al., (2022), we can use $g(O;\phi,c,d)$ as a moment function to estimate $\phi_{0}$. Evaluation of $g(O;\phi,c,d)$ replies on $\{\eta(x),p(z\mid u),p(y\mid u)\}$, which are directly targeted by the proposed approach. In particular, we note that the conditional outcome density among non-respondents can be expressed as
@@
which differs from (3) in terms of parameterization. When the baseline covariates include multiple continuous components, $\{\eta(x),p(z\mid u),p(y\mid u)\}$ can be estimated under user-specified, dimension-reducing parametric specifications $\{\eta(x;\xi),p(z\mid u;\beta),p(y\mid u;\psi)\}$, which allows for simpler conditions for asymptotic normality. Let $m(O;\gamma,\theta)$ denote the stacked moment functions in (2.5) for the parameter $\theta=(\xi^{\mathrm{\scriptscriptstyle T}},\beta^{\mathrm{\scriptscriptstyle T}},\psi^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. The proposed influence function-based estimator $(\hat{\phi}^{\mathrm{\scriptscriptstyle T}}(c,d),\hat{\theta}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ then solves
@@
Under standard regularity conditions for moment estimation (Newey and McFadden,, 1994), the estimator $\hat{\phi}(c,d)$ admits the following asymptotic expansion in the union model $\cup_{j=1,2}\mathcal{M}_{j}$,
$n^{1/2}(\hat{\phi}(c,d)-{\phi}_{0})=-n^{1/2}\left[E\biggr{\{}\frac{\partial}{\partial\phi}G(O;\phi,\bar{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\phi={\phi}_{0}}}\right]^{-1}\mathbb{P}_{n}\{G(O;{\phi}_{0},\bar{\theta},c,d)\}+o_{p}(1),$
@@
$-E\biggr{\{}\frac{\partial}{\partial\theta}g(O;{\phi}_{0},{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\theta=\bar{\theta}}}E\biggr{\{}\frac{\partial}{\partial\theta}m(O;{\gamma}_{0},{\theta})\biggr{\}}^{-1}{\biggr{\rvert}_{\theta=\bar{\theta}}}{m(O;\gamma,\bar{\theta})}.$
Furthermore, at the intersection submodel $\mathcal{M}_{1}\cap\mathcal{M}_{2}$ where all the working parametric models are correctly specified, $\hat{\phi}(c,d)$ admits the asymptotic expansion
@@
The first part of proposition 1 is due to the double robustness property that
$E\{{g}(O;\phi_{0},\bar{\xi},\bar{\beta},\bar{\psi},c,d)\}=0,$ (11)
if either $\{\eta(x;\bar{\xi}),p(z\mid u;\bar{\beta})\}=\{{\eta}(x),p(z\mid u)\}$ or $\{\eta(x;\bar{\xi}),p(y\mid u;\bar{\psi})\}=\{{\eta}(x),p(y\mid u)\}$. Thus, deviations of $p(z\mid u;\bar{\beta})$ or $p(y\mid u;\bar{\psi})$ away from the truth has no global effect on the moment condition. The second part of proposition 1 states that estimation of the nuisance parameter $\theta$ has no first-order effect on the asymptotic expansion of $\hat{\phi}(c,d)$ at the intersection submodel $\cap_{j=1,2}\mathcal{M}_{j}$ where $\{\eta(x;\bar{\xi}),p(z\mid u;\bar{\beta}),p(y\mid u;\bar{\psi})\}=\{{\eta}(x),p(z\mid u),p(y\mid u)\}$. This is a general property of estimators that are constructed based on influence functions due to Neyman orthogonality (Chernozhukov et al.,, 2022): varying $\theta$ away from $\bar{\theta}$ has no local effect on the moment condition, $E\{{\partial}g(O;\phi_{0},{\theta},c,d)/{\partial\theta}\}{\rvert_{\theta=\bar{\theta}}}=0$, which allows for simplification of the asymptotic variance formula. However, in general this simplification is lost as soon as one of the working models $p(z\mid u;{\beta})$ or $p(y\mid u;{\psi})$ is misspecified (Vermeulen and Vansteelandt,, 2015).
@@ 3.3 Local efficiency
When both $Y$ and $Z$ are discrete, the efficient influence function is indexed by the optimal choice $(c,d)=(c^{\ast},d^{\ast})$, which is characterized in the supplementary material. At the intersection submodel $\mathcal{M}_{1}\cap\mathcal{M}_{2}$, the influence function of $\hat{\phi}(c,d)$ coincides with $\Psi(O;\phi_{0},c,d)$, and hence the asymptotic variance of $\hat{\phi}(c^{\ast},d^{\ast})$ attains the semiparametric efficiency bound $\mathcal{V}=E\{\Psi(O;\phi_{0},c^{\ast},d^{\ast})\Psi(O;\phi_{0},c^{\ast},d^{\ast})^{\mathrm{\scriptscriptstyle T}}\}$ (local efficiency). In practice, we can implement a doubly robust and locally efficient estimator of $\phi_{0}$ based on a preliminary, consistent estimator of the optimal index functions, which we describe in the supplementary material. The results in Robins and Rotnitzky, (2001) show that the efficiency bound in the union model $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ coincides with the semiparametric efficiency bound $\mathcal{V}$. Therefore, $\hat{\phi}(c^{\ast},d^{\ast})$ also attains the efficiency bound in $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ at the intersection submodel $\mathcal{M}_{1}\cap\mathcal{M}_{2}$. When at least one of $Y$ and $Z$ is continuous, the efficient influence function is generally not available in closed form. In this case, we can construct a doubly robust and approximately locally efficient estimator of $\phi_{0}$, which is described in Sun et al., (2018).
@@ 3.4 A computationally simpler doubly robust estimator
The implementation of $\hat{\phi}(c,d)$ requires evaluation of the conditional outcome density among non-respondents given in (10). It is generally difficult to evaluate the integral in the denominator. For example, no closed form result for the logistic-normal integral is known, although various approximations have been proposed in the literature (Crouch and Spiegelman,, 1990). In this section, we propose a computationally simpler doubly robust estimator in the union model $\mathcal{M}_{1}\cup\mathcal{M}_{2}$ which avoids the numerical integration. The key observation is that double robustness (11) continues to to hold for the inverse propensity weighted function $\tilde{g}(O;\phi,c,d)={Rq(Y,X;\mu,c,d)}/{\pi(Y,X;\gamma)}$, although Neyman orthogonality in general no longer holds for this moment function due to the exclusion of the augmentation term involving non-respondents. Let $(\tilde{\phi}^{\mathrm{\scriptscriptstyle T}}(c,d),\hat{\theta}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ denote the joint solution to
@@
To illustrate the proposed computationally simpler doubly robust estimators, suppose $\mathcal{Y}=\mathcal{Z}=\{0,1\}$ and we set $c(U)=0$,
$\tilde{g}$ $(O;\phi,c=0,d)=\frac{R}{\pi(Y,X;\gamma)}\left[\begin{matrix}Y-\mu\\ d_{yz}(U)\{Y-p(Y=1\mid U)\}\{Z-p(Z=1\mid U)\}\end{matrix}\right],$
for some $d_{yz}(u)\in\mathbb{R}^{p_{\gamma}}$. In this case, following the examples in section 2.5, estimation can proceed in two stages. In the first stage, $(\tilde{\gamma}^{\mathrm{\scriptscriptstyle T}},\tilde{\theta}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ solves
$\mathbb{P}_{n}\left[\frac{R}{\pi(Y,X;\xi,\gamma)}d_{yz}(U)\{Y-p(Y=1\mid U;\psi)\}\{Z-p(Z=1\mid U;\beta)\}\right]=0,$
jointly with (2.5). Then the Hájek, (1971) estimator $\tilde{\mu}$ of the outcome mean can be constructed with the estimated inverse propensity weights $\{R_{i}/\pi(Y_{i},X_{i};\tilde{\xi},\tilde{\gamma}):i=1,...,n\}$.
@@
Under standard regularity conditions for moment estimation (Newey and McFadden,, 1994), the estimator $\tilde{\phi}(c,d)$ admits the following asymptotic expansion in the union semiparametric model $\mathcal{M}_{1}\cup\mathcal{M}_{2}$,
$n^{1/2}(\tilde{\phi}(c,d)-{\phi}_{0})=-n^{1/2}\left[E\biggr{\{}\frac{\partial}{\partial\phi}\tilde{G}(O;\phi,\bar{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\phi={\phi}_{0}}}\right]^{-1}\mathbb{P}_{n}\{\tilde{G}(O;{\phi}_{0},\bar{\theta},c,d)\}+o_{p}(1),$
@@
$-E\biggr{\{}\frac{\partial}{\partial\theta}\tilde{g}(O;{\phi}_{0},{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\theta=\bar{\theta}}}E\biggr{\{}\frac{\partial}{\partial\theta}m(O;{\gamma}_{0},{\theta})\biggr{\}}^{-1}{\biggr{\rvert}_{\theta=\bar{\theta}}}{m(O;\gamma,\bar{\theta})}.$
The asymptotic variance of $\tilde{\phi}(c,d)$ cannot attain the efficiency bound in $\mathcal{M}$ (and also in $\mathcal{M}_{1}\cup\mathcal{M}_{2}$) at the intersection submodel $\mathcal{M}_{1}\cap\mathcal{M}_{2}$, even with the optimal choice of index functions in the discrete case. Therefore, the estimator $\tilde{\phi}(c,d)$ is doubly robust but not locally efficient, and intuitively this is because it fails to incorporate the information from non-respondents when all the working models are correctly specified. Nonetheless, because we have already paid homage to the need for efficiency by using parametric models, the potential prize of attempting to attain local efficiency may not always be worth the chase in view of the additional computational demands. As the goal of this paper is to produce a statistically sound and practically useful method to tackle nonignorable missing data, we will focus on the estimator $\tilde{\phi}(c,d)$ in the simulation study and application.
@@ 4 Simulation Studies
In order to investigate the finite-sample properties of the doubly robust estimator proposed in Section 3, we perform Monte Carlo simulations involving identical and independently generated data $\{O_{1},...,O_{n}\}$. The baseline covariate $U=(U_{1},U_{2})^{\mathrm{\scriptscriptstyle T}}$ is generated from a bivariate normal distribution $N(0,\Sigma)$, where the elements of $\Sigma$ are $\sigma^{2}_{1}=\sigma^{2}_{2}=1$ and $\sigma_{12}=0.2$. Conditional on $U$, $(R,Y,Z)$ is generated from the following generalized linear models (GLMs) consistent with assumptions 13,
@@
where $\gamma=2$. We implement the proposed doubly robust estimator $\tilde{\phi}(c,d)=(\tilde{\mu},\tilde{\gamma})^{\mathrm{\scriptscriptstyle T}}$ with $c(y,x)=0$, $d(y,x)=yz$ based on the GLMs $\pi(y,x;\xi,\gamma)=\textup{expit}\{(h^{\mathrm{\scriptscriptstyle T}}_{1}(u),z)^{\mathrm{\scriptscriptstyle T}}\xi+\gamma y\}$, $p(Z=1\mid u;\beta)=\textup{expit}\{h_{2}(u)\beta\}$ and $p(Y=1\mid u;\psi)=\textup{expit}\{h_{3}(u)\psi\}$ under the following five scenarios. The model for $\eta(x)$, $p(Z=1\mid u)$ or $p(Y=1\mid u)$ is misspecified if $h_{j}(u)=(1,u_{1},u^{2}_{1})$ for $j=1$, 2 or 3, respectively.
@@
We compare $\tilde{\mu}$ with the doubly robust estimator $\hat{\mu}_{{\textup{dr}}}$ proposed by Sun et al., (2018), and specify the GLM $p(Y=1\mid R=1,x;\psi_{1},\lambda)=\textup{expit}\{h_{3}(u)\psi_{1}+\lambda z\}$ under (C1)–(C5) to evaluate its use in practice. In addition, we implement the complete-case estimator of the outcome mean $\hat{\mu}_{{\textup{cc}}}=\mathbb{P}_{n}(RY)$, as well as the infeasible full-data estimator $\hat{\mu}_{{\text{full}}}=n^{-1}\sum^{n}_{i=1}Y_{i}$ as performance benchmark. For inference, we construct 95% Wald confidence intervals based on the sandwich estimator of asymptotic variance.
The following remarks can be made based on the results of 1000 simulation replicates of sample size $n=500$, $1000$ or $5000$ summarized in Table 1. The complete-case estimator $\hat{\mu}_{{\textup{cc}}}$ exhibits severe bias and undercoverage. In agreement with theory, $\tilde{\mu}$ performs well in terms of bias and coverage cross scenarios (C1)–(C3), but is biased under (C4) and (C5) with misspecified model for $\eta(x)$. The estimator $\hat{\mu}_{{\textup{dr}}}$ shows negligible bias and coverage proportion close to the nominal level under scenarios (C1) and (C3), but is biased under (C2) and (C5) with misspecified model for $p(Z=1\mid u)$. It also has small but noticeable bias under (C4). The relative efficiency of $\tilde{\mu}$ compared to the full-data estimator $\hat{\mu}_{{\textup{full}}}$ is approximately $0.15$ based on Monte Carlo variance at $n=5000$. The supplementary material contains additional Monte Carlo simulation results with a continuous $Y$.

### v4

@@ 3.2 Influence function-based doubly robust estimators
Following Robins and Rotnitzky (2001) and Chernozhukov et al. (2022), we can use $g(O;\phi,c,d)$ as a moment function to estimate $\phi_{0}$. Evaluation of $g(O;\phi,c,d)$ relies on $\{\eta(x),p(z\mid u),p(y\mid u)\}$, which are directly targeted by the proposed approach. In particular, we note that the conditional outcome density among non-respondents can be expressed as
@@
which differs from (2.4) in terms of parameterization. When the baseline covariates include multiple continuous components, $\{\eta(x),p(z\mid u),p(y\mid u)\}$ can be estimated under user-specified, dimension-reducing parametric specifications $\{\eta(x;\xi),p(z\mid u;\beta),p(y\mid u;\psi)\}$, which allows for simpler conditions for asymptotic normality. Let $m(O;\gamma,\theta)$ denote the stacked moment functions in (2.5) for the parameter $\theta=(\xi^{\mathrm{\scriptscriptstyle T}},\beta^{\mathrm{\scriptscriptstyle T}},\psi^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$. The proposed influence function-based estimator $(\hat{\phi}^{\mathrm{\scriptscriptstyle T}}(c,d),\hat{\theta}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ then solves
@@
Under standard regularity conditions for moment estimation (Newey and McFadden, 1994), the estimator $\hat{\phi}(c,d)$ admits the following asymptotic expansion in the union model $\cup_{j=1,2}\mathcal{M}_{j}$,
$n^{1/2}(\hat{\phi}(c,d)-{\phi}_{0})=-n^{1/2}\left[\mathbb{E}\biggr{\{}\frac{\partial}{\partial\phi}G(O;\phi,\bar{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\phi={\phi}_{0}}}\right]^{-1}\mathbb{P}_{n}\{G(O;{\phi}_{0},\bar{\theta},c,d)\}+o_{p}(1),$
@@
$-\mathbb{E}\biggr{\{}\frac{\partial}{\partial\theta}g(O;{\phi}_{0},{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\theta=\bar{\theta}}}\mathbb{E}\biggr{\{}\frac{\partial}{\partial\theta}m(O;{\gamma}_{0},{\theta})\biggr{\}}^{-1}{\biggr{\rvert}_{\theta=\bar{\theta}}}{m(O;\gamma,\bar{\theta})}.$
Furthermore, at the intersection submodel $\cap_{j=1,2}\mathcal{M}_{j}$ where all the working parametric models are correctly specified, $\hat{\phi}(c,d)$ admits the asymptotic expansion
@@
The first part of proposition 1 is due to the following double robustness property,
$\mathbb{E}\{{g}(O;\phi_{0},\bar{\xi},\bar{\beta},\bar{\psi},c,d)\}=0,$ (3.12)
if either $\{\eta(x;\bar{\xi}),p(z\mid u;\bar{\beta})\}=\{{\eta}(x),p(z\mid u)\}$ or $\{\eta(x;\bar{\xi}),p(y\mid u;\bar{\psi})\}=\{{\eta}(x),p(y\mid u)\}$. Thus, deviations of $p(z\mid u;\bar{\beta})$ or $p(y\mid u;\bar{\psi})$ away from the truth has no global effect on the moment condition. The second part of proposition 1 states that estimation of the nuisance parameter $\theta$ has no first-order effect on the asymptotic expansion of $\hat{\phi}(c,d)$ at the intersection submodel $\cap_{j=1,2}\mathcal{M}_{j}$ where $\{\eta(x;\bar{\xi}),p(z\mid u;\bar{\beta}),p(y\mid u;\bar{\psi})\}=\{{\eta}(x),p(z\mid u),p(y\mid u)\}$. This is a general property of estimators that are constructed based on influence functions due to Neyman orthogonality (Chernozhukov et al., 2022), which allows for simplification of the asymptotic variance formula. However, in general this simplification is lost as soon as one of the working models $p(z\mid u;{\beta})$ or $p(y\mid u;{\psi})$ is misspecified (Vermeulen and Vansteelandt, 2015).
@@ 3.3 Local efficiency
When both $Y$ and $Z$ are discrete, the efficient influence function in $\mathcal{M}$ is indexed by the optimal choice $(c,d)=(c^{\ast},d^{\ast})$, which is characterized in the supplementary material. At the intersection submodel $\cap_{j=1,2}\mathcal{M}_{j}$, the influence function of $\hat{\phi}(c,d)$ coincides with $\Psi(O;\phi_{0},c,d)$, and hence the asymptotic variance of $\hat{\phi}(c^{\ast},d^{\ast})$ attains the semiparametric efficiency bound $\mathcal{V}=\mathbb{E}\{\Psi(O;\phi_{0},c^{\ast},d^{\ast})\Psi(O;\phi_{0},c^{\ast},d^{\ast})^{\mathrm{\scriptscriptstyle T}}\}$ (local efficiency). In practice, we can implement a doubly robust and locally efficient estimator of $\phi_{0}$ based on a preliminary, consistent estimator of the optimal index functions, which we describe in the supplementary material. The results in Robins and Rotnitzky (2001) show that the efficiency bound in the union model $\cup_{j=1,2}\mathcal{M}_{j}$ coincides with the semiparametric efficiency bound $\mathcal{V}$ in $\mathcal{M}$. Therefore, $\hat{\phi}(c^{\ast},d^{\ast})$ also attains the efficiency bound in $\cup_{j=1,2}\mathcal{M}_{j}$ at the intersection submodel $\cap_{j=1,2}\mathcal{M}_{j}$. When at least one of $Y$ and $Z$ is continuous, the efficient influence function is typically not available in closed form. In this case, we can construct a doubly robust and approximately locally efficient estimator of $\phi_{0}$, which is described in Sun et al. (2018).
@@ 3.4 A computationally simpler doubly robust estimator
The implementation of $\hat{\phi}(c,d)$ requires evaluation of the conditional outcome density among non-respondents given in (3.11). It is generally difficult to evaluate the integral in the denominator. For example, no closed form result for the logistic-normal integral is known, although various approximations have been proposed in the literature (Crouch and Spiegelman, 1990). In this section, we propose a computationally simpler doubly robust estimator in the union model $\cup_{j=1,2}\mathcal{M}_{j}$ which avoids the numerical integration. The key observation is that double robustness (3.12) continues to hold for the inverse propensity weighted function $\tilde{g}(O;\phi,c,d)={Rq(W;\mu,c,d)}/{\pi(W;\gamma)}$ which excludes the augmentation term involving non-respondents. Let $(\tilde{\phi}^{\mathrm{\scriptscriptstyle T}}(c,d),\hat{\theta}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ denote the joint solution to
@@
To illustrate the proposed computationally simpler doubly robust estimators, suppose $\mathcal{Y}=\mathcal{Z}=\{0,1\}$ and we set $c(W)=0$. In this case, following the examples in section 2.5, estimation based on $\tilde{g}(O;\phi,c=0,d)$ can proceed in two stages. In the first stage, $(\tilde{\gamma}^{\mathrm{\scriptscriptstyle T}},\tilde{\theta}^{\mathrm{\scriptscriptstyle T}})^{\mathrm{\scriptscriptstyle T}}$ solves (2.5) jointly with
$\mathbb{P}_{n}\left[\frac{R}{\pi(W;\xi,\gamma)}d_{yz}(U)\{Y-p(Y=1\mid U;\psi)\}\{Z-p(Z=1\mid U;\beta)\}\right]=0,$
for some $d_{yz}(U)\in\mathbb{R}^{p_{\gamma}}$. Then the Hájek (1971) estimator $\tilde{\mu}$ of the outcome mean can be constructed with the estimated inverse propensity weights $\{R_{i}/\pi(W_{i};\tilde{\xi},\tilde{\gamma}):i=1,...,n\}$.
@@
Under standard regularity conditions for moment estimation (Newey and McFadden, 1994), the estimator $\tilde{\phi}(c,d)$ admits the following asymptotic expansion in the union model $\cup_{j=1,2}\mathcal{M}_{j}$,
$n^{1/2}(\tilde{\phi}(c,d)-{\phi}_{0})=-n^{1/2}\left[\mathbb{E}\biggr{\{}\frac{\partial}{\partial\phi}\tilde{G}(O;\phi,\bar{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\phi={\phi}_{0}}}\right]^{-1}\mathbb{P}_{n}\{\tilde{G}(O;{\phi}_{0},\bar{\theta},c,d)\}+o_{p}(1),$
@@
$-\mathbb{E}\biggr{\{}\frac{\partial}{\partial\theta}\tilde{g}(O;{\phi}_{0},{\theta},c,d)\biggr{\}}{\biggr{\rvert}_{\theta=\bar{\theta}}}\mathbb{E}\biggr{\{}\frac{\partial}{\partial\theta}m(O;{\gamma}_{0},{\theta})\biggr{\}}^{-1}{\biggr{\rvert}_{\theta=\bar{\theta}}}{m(O;\gamma,\bar{\theta})}.$
The asymptotic variance of $\tilde{\phi}(c,d)$ cannot attain the efficiency bound in $\mathcal{M}$ (and hence also in $\cup_{j=1,2}\mathcal{M}_{j}$) even at the intersection submodel $\cap_{j=1,2}\mathcal{M}_{j}$. Therefore, the estimator $\tilde{\phi}(c,d)$ is doubly robust but not locally efficient, and intuitively this is because it fails to incorporate the information from non-respondents when all the working models are correctly specified. Nonetheless, because we have already paid homage to the need for efficiency by using parametric models, the potential prize of attempting to attain local efficiency may not always be worth the chase in view of the additional computational demands. As the goal of this paper is to produce a statistically sound and practically useful method to tackle nonignorable missing data, we will focus on the estimator $\tilde{\phi}(c,d)$ in the simulation study and application.
@@ 4 Simulation Studies
In order to investigate the finite-sample properties of the doubly robust estimator proposed in Section 3, we perform Monte Carlo simulations involving identical and independently generated data $\{O_{1},...,O_{n}\}$. The baseline covariate $U=(U_{1},U_{2})^{\mathrm{\scriptscriptstyle T}}$ is generated from a bivariate normal distribution $N(0,\Sigma)$, where the elements of $\Sigma$ are $\sigma^{2}_{1}=\sigma^{2}_{2}=1$ and $\sigma_{12}=0.2$. Conditional on $U$, $(R,Y,Z)$ is generated from the following generalized linear models consistent with assumptions 13,
@@
where $\gamma=2$. We implement the proposed doubly robust estimator $\tilde{\phi}(c,d)=(\tilde{\mu},\tilde{\gamma})^{\mathrm{\scriptscriptstyle T}}$ with $c(w)=0$, $d(w)=yz$ based on the models $\pi(w;\xi,\gamma)=\textup{expit}\{(h^{\mathrm{\scriptscriptstyle T}}_{1}(u),z)^{\mathrm{\scriptscriptstyle T}}\xi+\gamma y\}$, $p(Z=1\mid u;\beta)=\textup{expit}\{h_{2}(u)\beta\}$ and $p(Y=1\mid u;\psi)=\textup{expit}\{h_{3}(u)\psi\}$ under the following five scenarios. The model for $\eta(x)$, $p(Z=1\mid u)$ or $p(Y=1\mid u)$ is misspecified if $h_{j}(u)=(1,u_{1},u^{2}_{1})$ for $j=1$, 2 or 3, respectively.
@@
To compare $\tilde{\mu}$ with the doubly robust estimator $\hat{\mu}_{{\textup{dr}}}$ proposed by Sun et al. (2018), we specify the working parametric model $p(Y=1\mid R=1,x;\psi_{1},\lambda)=\textup{expit}\{h_{3}(u)\psi_{1}+\lambda z\}$ under (C1)–(C5) to evaluate its use in practice. In addition, we implement the complete-case estimator of the outcome mean $\hat{\mu}_{{\textup{cc}}}=\mathbb{P}_{n}(RY)$, as well as the infeasible full-data estimator $\hat{\mu}_{{\text{full}}}=n^{-1}\sum^{n}_{i=1}Y_{i}$ as performance benchmark. For inference, we construct 95% Wald confidence intervals based on the sandwich estimator of asymptotic variance.
The following remarks can be made based on the results of 1000 simulation replicates of sample size $n=500$, $1000$ or $5000$ summarized in Table 1. The complete-case estimator $\hat{\mu}_{{\textup{cc}}}$ exhibits severe bias and undercoverage. In agreement with theory, $\tilde{\mu}$ performs well in terms of bias and coverage cross scenarios (C1)–(C3), but is biased under (C4) and (C5) with misspecified model for $\eta(x)$. The estimator $\hat{\mu}_{{\textup{dr}}}$ shows negligible bias and coverage proportion close to the nominal level under scenarios (C1) and (C3), but is biased under (C2) and (C5) with misspecified model for $p(Z=1\mid u)$. It also has small but noticeable bias under (C4). The relative efficiency of $\tilde{\mu}$ compared to the full-data estimator $\hat{\mu}_{{\textup{full}}}$ is approximately $0.15$ based on Monte Carlo variance at $n=5000$. The supplementary material contains additional Monte Carlo simulation results under violations of the exclusion restriction in assumption 2, as well as with a continuous $Y$.